OpenStudy (anonymous):

MEDALS AWARDED!!! Determine whether the vectors u and v are parallel, orthogonal, or neither. u = <1, -2>, v = <-4, 8>

OpenStudy (anonymous):

is it orthdontal?

OpenStudy (anonymous):

@wilfley !!! good to see you :)

OpenStudy (anonymous):

hey

OpenStudy (anonymous):

did i answer correctly?

OpenStudy (anonymous):

if two vectors are perpendicular (orthogonal), they will have a dot product of 0 i think...let me check

OpenStudy (anonymous):

my answer choices are : Orthogonal Neither Parallel

OpenStudy (anonymous):

yeah, they aren't orthogonal because the dot product is -20 which of course isn't 0 lol

OpenStudy (anonymous):

they are parallel is the cross product of the two vectors is 0, lets see.

OpenStudy (anonymous):

sso it would be parrallel??

OpenStudy (anonymous):

\[\left[\begin{matrix}1 & -2 \\ -4 & 8\end{matrix}\right]\] (1*8)-[-4*(-2)] =0 so yeah, parallel

OpenStudy (anonymous):

thank you! and one more

OpenStudy (anonymous):

Write the complex number in the form a + bi. 3(cos 60° + i sin 60°)

OpenStudy (anonymous):

is it 3/2+3sqrtof3/2i ???

OpenStudy (anonymous):

crap, I don't remember that one, what is the chapter about?

OpenStudy (anonymous):

im not sure.. i cant leave the test

OpenStudy (anonymous):

is the i bold for a vector or regular for imaginary numbers?